OpenStudy (anonymous):

How do I find the zero of this function? (click to see equation, I'm working on it)

OpenStudy (anonymous):

\[F^-1(x)\] if \[F(x) = x + 3\]

OpenStudy (anonymous):

If \[ F(x)=x+3 \] then \[ F^{-1}(x) = x-3 \] Can you find the zero of that?

OpenStudy (anonymous):

It is zero when x = 3

OpenStudy (anonymous):

\[f(0)=0+3=3....................f ^{-1}=0-3=-3\]

OpenStudy (zarkon):

\[F(0)=3\] so \[F^{-1}(3)=0\]

OpenStudy (anonymous):

I think I get it, but what does the -1 on the F mean? what am I supposed to do to the equation? I thought a negative exponent turned it into a fraction...?

OpenStudy (zarkon):

inverse function

OpenStudy (zarkon):

YTheManifold gave you the inverse function if you want it

OpenStudy (anonymous):

Ohh, now I get it. Thank you guys! :)